OpenStudy (studytheworld):

Help!!! Will medal and fan! True or false? The distance between points (1,2) and (x1,y1) is the square root of (x1-1)^2+(y1-2)^2.

OpenStudy (danjs):

yeah for a line connecting 2 points you can form a right triangle with the legs as the components in x and y direction lengths d^2 = side^2 + side2^2 pythagorean a^2+b^2=c^2 for right triangle lengths

OpenStudy (danjs):

there it is 

OpenStudy (studytheworld):

Okay, true. Thanks :)